OpenStudy (anonymous):

Find the volume of a pyramid which has a square base of 12 inches and a height of 24 inches.864? A right circular cylinder has all of its dimensions multiplied by three. How is its surface area affected? by 9?

OpenStudy (lgbasallote):

by height you mean vertical height right? if so you can use the formula \[\large V = \frac{1}{3}(base)(height)\]

OpenStudy (lgbasallote):

base means base area btw

OpenStudy (anonymous):

its 96 na???????? bcos 1/3*base*height= 1/3*12*24=96

OpenStudy (anonymous):

okso 96 for number 1

OpenStudy (anonymous):

nuumber 2 is correct?

OpenStudy (lgbasallote):

if dimensions mean radius and height then 9 is right

OpenStudy (anonymous):

2nd is ri8

OpenStudy (anonymous):

ok thx